PHOTODISSOCIATION DYNAMICS OF CHLORINE DIOXIDE

\underline{THOMAS J. DOWD}, LEANNE M. MILLER, and JAMES S. KELLER, Department of Chemistry and Biochemistry, University of Notre Dame, Notre Dame IN, 46556.

A ground state population depletion grating was used to elucidate the effects of initial nuclear motion on the photodissociation dynamics of chlorine dioxide. The ultraviolet absorption of OClO occurs via the A^{2}A_{2} ← X^{2}B_{1} transition, but dissociation of OClO* proceeds on several different excited state surfaces leading to two distinct product channels. Since several excited bending states (ν_{2}) are populated at room temperature, a depletion grating can be formed involving these vibrational states by the interference of two degenerate pump beams tuned to transitions which lie below the predissociation threshold. A probe beam was tuned through the dissociation region to record the A ← X absorption spectra of specific vibrational states.